### v1

@@ 3.3 Optimal hypersurface decision tree algorithm
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: number of splitting rules; $G$: dimension of the embedded space
  2. 2.
    Output: Array of $\left(k,D\right)$-nested-combinations for $k\in\left\{0,\ldots K\right\}$
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    $\rho_{M}\left(\mathit{xs}\right)=\mathit{embed}\left(M,\mathit{xs}\right)$ // map the data list $\mathit{xs}$ to the embedded space
  7. 7.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  8. 8.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  9. 9.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  10. 10.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  11. 11.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  12. 12.
    $\mathit{css}\left[G\right]=\left[\>\right]$
  13. 13.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  14. 14.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  15. 15.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  16. 16.
        $\mathit{ncss}\left[k\right]=\mathit{updates}_{\text{arMat}}\left(i,\mathit{ncss}\left[k\right],\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$
  17. 17.
@@
  18. 18.
    if $E_{\text{0-1}}\left(t_{\text{best}}\right)\leq E_{\text{0-1}}\left(t_{\text{opt}}\right)$:
  19. 19.
      $t_{\text{opt}}=t_{\text{best}}$
  20. 20.
    $\mathit{ncss}\left[K\right]=\left[\>\right]$ // eliminate size $K$ nested combinations after evaluation
  21. 21.
    return $t_{\text{opt}}$
Algorithm 2 $\mathit{hodt}\left(K,M,\mathit{xs}\right)$
@@
The main difference between the brute-force specification 5 and Algorithm 3 lies in the order of execution. In the specification, functions are composed sequentially, meaning each function begins only after the previous one has completed. If written as pseudocode, both $\mathit{min}{}_{E_{\text{0-1}}}$ and $\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}$ would appear outside the outermost loop that defines $\mathit{\mathit{nestedCombs}}\left(K,G\right)$. This is inefficient because $\mathit{\mathit{nestedCombs}}\left(K,G\right)$ generates a prohibitively large number of candidate solutions.
@@ 4 Heuerestic methods
  1. 1.
    Input: Parameteres for coreset: $\mathit{BS}$: Block size; $R$: number of shuffle time in each filtering process; $L$: Max-heap size; $B_{\max}$: Maximum input size for the Deep-ICE algorithm; $c\in\left(0,1\right]$: Shrinking factor for heap size
  2. 2.
      Parameteres for $\mathit{hodt}$: $K$: number of splitting rules, $M$ degree of polynomial hypersurface splitting rules; $\mathit{xs}$: input data list;
  3. 3.
    Output: Max-heap containing top $L$ configurations and associated data blocks
  4. 4.
    Initialize coreset $\mathcal{C}\leftarrow\mathit{ds}$
  5. 5.
    while $\mathcal{C}\leq B_{\max}$ do:
  6. 6.
    Reshuffle the data, divide $\mathcal{C}$ into $\left\lceil\frac{\left|\mathcal{C}\right|}{\mathit{BS}}\right\rceil$ blocks $\mathcal{C}_{B}=\left\{C_{1},C_{2},\ldots,C_{\left\lceil\frac{\left|\mathcal{C}\right|}{\mathit{BS}}\right\rceil}\right\}$
  7. 7.
    Initialize a size $L$ max-heap $\mathcal{H}_{L}$
  8. 8.
    for $r\leftarrow 1$ to $R$ do:
  9. 9.
      $r=r+1$
  10. 10.
      for $C\in\mathcal{C}_{B}$ do:
  11. 11.
       $\mathit{cnfg}\leftarrow\mathit{hodt}\left(K,M,C\right)$
  12. 12.
       $\mathcal{H}_{L}.\text{push}\left(\mathit{cnfg},C\right)$
  13. 13.
    $\mathcal{C}\leftarrow\mathit{unique}\left(\mathcal{H}_{L}\right)$ // Merge blocks and remove duplicates
  14. 14.
    $L\leftarrow L\times c$ // Shrink heap size:
  15. 15.
    $\mathit{cnfg}\leftarrow\mathit{hodt}\left(K,M,\mathcal{C}\right)$ // Final refinement
  16. 16.
    $\mathcal{H}_{L}.\text{push}\left(\mathit{cnfg},\mathcal{C}\right)$
  17. 17.
    return $\mathcal{H}_{L}$
Algorithm 3 $\mathit{hodt}Coreset\left(K,M,\mathit{xs},\mathit{BS},R,L,B_{\max},c\right)$
@@ 4.2 $\mathit{sodt}$ with selected hyperplanes
  1. 1.
    Input: $K$: number of splitting rules, $M$ degree of polynomial hypersurface splitting rules; $\alpha$: threshold for duplicate data $\mathit{xs}$: input data list;
  2. 2.
    Output: The updated nested combinations $\mathit{NCs}$ without crossed hyperplanes, a vector of e $\left(k+2\right)\times\left(k+1\right)$ matrices
  3. 3.
    $t_{\text{best}}=\mathit{empty}\left(K,1\right)$ // a vector
  4. 4.
    $\mathcal{H}_{C}=\mathit{\mathit{hodt}Coreset}\left(1,M,\mathit{xs},BS,R,L,B_{\max},c\right)$ // selecting candidate hypersurfaces with low 0-1 loss
  5. 5.
    $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$
  6. 6.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  7. 7.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  8. 8.
    for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,n+1\right)\right)$ do:
  9. 9.
      $\mathit{ncss}^{\prime}=\mathit{filter}_{q\left(\alpha\right)}\left(\mathit{map}\left(\cup\left[n\right],\mathit{ncss}\left[k-1\right]\right)\right)$ // filter out nested combinations $\mathit{nc}$ in $\mathit{ncss}^{\prime}\left[k\right]$ with duplicates data points small than threshold $\alpha$
  10. 10.
      $t_{\text{best}}\left[k\right]=\mathit{min}_{E_{\text{0-1}}}\left(\mathit{mapL}\left(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}^{\prime}\right)\right)$
  11. 11.
      $\mathit{ncss}\left[k\right]=\mathit{ncss}\left[k\right]\cup\mathit{reverse}\left(\mathit{ncss}^{\prime}\right)$ // update $\mathit{ncss}$ to generate combinations in revolving door ordering,
  12. 12.
    $\mathit{ncss}\left[K\right]=\left[\>\right]$
  13. 13.
    return $t_{\text{best}}$
Algorithm 4 $\mathit{sodtWSH}\left(K,M,\alpha,\mathit{xs}\right)$
@@ 5.1 Computational analysis
In this section, we analyze in detail the computational efficiency of $\mathit{sodt}$ for solving the size-constrained ODT problem. Since $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ share similar advantages—both are fully vectorizable but unable to exploit efficient dynamic programming (DP)—we focus on comparing $\mathit{sodt}_{\text{vec}}$ against the dynamic programming algorithm $\mathit{sodt}_{\text{rec}}$.
@@ 5.1.1 Comparison between the vectorized and recursive implementation
Figure 5: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on sequential setting. .
Figure 6: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on parallel setting..
@@
Following the detailed comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ in both sequential and parallel settings, we now turn to the computational scalability of the winning method, $\mathit{sodt}_{\text{vec}}$. We focus on analyzing the pure computational scalability of the algorithm by benchmarking its performance without any acceleration techniques. Specifically, we measure the total wall-clock runtime of $\mathit{sodt}_{\text{vec}}$ for solving $1\times 10^{6}$ instances of feasible nested combinations (with no crossed hyperplanes) for different values of KK。In our setup, cases with $K\leq 4$ are executed on the CPU, while cases with $K>4$ are executed on the GPU.
@@ 5.2 Combinatorial complexity of the hyperplane/hypersurface decision tree
5.2.1 Combinatorial complexity of nested combinations after fitlered out crossed hyperplanes
Figure 8: Combinatorial complexity of feasible nested combinations with varying $K$. Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) (leftmost bar) to the value of $K$ labeled below each group of bars.
@@ 5.2.1 Combinatorial complexity of nested combinations after fitlered out crossed hyperplanes
Comparing the combinatorial complexity of hypersurface and hyperplanes in the same diemension
Figure 9: Combinatorial complexity of feasible nested combinations with varying $K$. Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) (leftmost bar) to the value of $K$ labeled below each group of bars.
@@ 5.3 Computational experiments with synthetic datasets
2
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.12 (0.04)
93.82 (3.01)
91.72 (3.83)
100 (0.00)
100 (0.00)
82.61 (0.04)
86.12 (3.73)
84.57 (4.03)
93.50 (3.21)
100 (0.00)
2.92 (0.27)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.49 (0.16)
2.00 (0.00)
1.67 (1.20)
1.67 (1.20)
1.96 (0.15)
2.15 (0.20)
2.00 (0.00)
1.65 (0.16)
1.66 (0.16)
3
CART-depth-3
CART-size-6
ConTree-3
HODT
Ground Truth
90.72 (047)
93.88 (3.89)
95.98 (2.93)
100 (0.00)
100 (0.00)
81.30 (5.71)
83.92 (5.55)
85.25 (4.96)
94.91 (2.36)
1 (0.00)
5.90 (0.00)
6.00 (0.00)
6.94 (0.24)
3.00 (0.00)
3.00 (0.00)
3.00 (0.00)
4.02 (0.65)
3.00 (0.00)
3.00 (0.00)
2.72 (0.45)
2.83 (0.14)
3.10 (0.24)
2.99 (0.03)
2.25 (0.0)
2.18 (0.11)
2.77 (0.19)
2.71 (0.20)
2.99 (0.04)
2.18 (0.25)
2.07 (0.24)
4
CART-depth-3
CART-size-8
ConTree-3
HODT
Ground Truth
87.98 (5.26)
93.90 (3.93)
93.12 (4.02)
97.46 (1.70)
100 (0.00)
77.66 (5.38)
81.93 (4.93)
81.51 (4.78)
89.89 (3.71)
100 (0.00)
6.12 (90.86)
8.00 (0.00)
7.00 (0.00)
4.00 (0.00)
4.00 (0.00)
3.00 (0.00)
4.54 (0.57)
3.00 (0.00)
3.52 (0.50)
3.52 (0.50)
2.87 (0.14)
3.49 (0.21)
3.00 (0.00)
2.63 (0.19)
2.63 (0.19)
2.86 (0.15)
3.13 (0.21)
3.00 (0.00)
2.60 (0.31)
2.63 (0.33)
@@
100
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.76 (5.11)
97.62 (2.51)
93.52 (3.53)
100 (0.00)
100 (0.00)
75.46 (15.70)
80.26 (16.80)
78.62 (16.14)
84.38 (19.27)
100 (0.00)
2.88 (0.32)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
0.00 (0.00)
2.00 (0.00)
3.92 (0.69)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.96 (0.11)
3.01 (0.30)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.96 (0.12)
2.52 (0.33)
2.00 (0.00)
1.66 (0.25)
1.70 (0.25)
200
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.00 (4.49)
94.49 (3.02)
92.87 (3.38)
99.94 (0.16)
100 (0.00)
84.62 (5.23)
89.57 (4.04)
88.51 (3.99)
97.97 (1.01)
100 (0.00)
2.92.00 (0.27)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.26 (0.44)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.50 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.98 (0.07)
2.17 (0.21)
2.00 (0.00)
1.71 (0.25)
1.70 (0.25)
400
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
87.45 (4.69)
92.58 (3.48)
91.44 (3.37)
99.85 (0.21)
100 (0.00)
84.62 (4.90)
89.35 (4.08)
88.78 (4.00)
98.98 (0.58)
100 (0.00)
2.98 (0.14)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
2.53 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.99 (0.05)
2.15 (0.24)
2.00 (0.00)
1.71 (0.25)
1.70 (0.25)
800
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
86.91 (4.63)
92.21 (3.43)
91.00 (3.59)
99.87 (0.14)
100 (0.00)
85.14 (4.96)
90.12 (3.82)
89.18 (3.94)
99.40 (0.36)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.3 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.53 (0.18)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.13 (0.22)
2.00 (0.00)
1.69 (0.25)
1.70 (0.25)
1600
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
85.66 (5.05)
91.48 (3.69)
90.13 (3.69)
99.87 (0.12)
100 (0.00)
84.97 (5.15)
90.53 (3.71)
89.38 (3.84)
99.65 (0.22)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.51 (0.16)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.16 (0.26)
2.00 (0.00)
1.70 (0.24)
1.70 (0.25)

### v2

@@ 3.3 Optimal hypersurface decision tree algorithm
Algorithm 2 OHDT with Thinning ($hodt(K,M,xs)$)
1:Input: $\mathit{xs}$: data list, length NN; $K$: splitting rules; $G$: embedded space dimension
2:Output: Array of $(k,D)$-nested-combinations for $k\in\{0,\ldots,K\}$
@@
[…]
7:for $n\in\mathit{range}(0,N)$ do
8:  for $j\in\mathit{reverse}(\mathit{range}(G,n+1))$ do
9:   $\mathit{updates}\leftarrow\mathit{reverse}(\mathit{map}(\cup\rho_{M}(\mathit{xs})[n],\mathit{css}[j-1]))$
10:   $\mathit{css}[j]\leftarrow\mathit{css}[j]\cup\mathit{updates}$ $\triangleright$ Update combinations
11:   $\mathit{asgn}^{+},\mathit{asgn}^{-}\leftarrow\mathit{genModels}(\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Generate predictions
12:   $\mathit{css}[G]\leftarrow[\,]$
13:  end for
14:  $C_{1}\leftarrow\binom{n}{G}$, $C_{2}\leftarrow\binom{n}{G-1}$
15:  for $i\in\mathit{range}(C_{1},C_{2})$ do
16:   for $k\in\mathit{reverse}(\mathit{range}(K,i+1))$ do
17:    $\mathit{ncss}[k]\leftarrow\mathit{updates}_{arMat}(i,\mathit{ncss}[k],\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$
18:   end for
19:  end for
20:  $t_{best}\leftarrow\min_{E_{0-1}}(\mathit{mapL}(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}[K]))$ $\triangleright$ Optimal decision tree
21:  if $E_{0-1}(t_{best})\leq E_{0-1}(t_{opt})$ then
22:   $t_{opt}\leftarrow t_{best}$
23:  end if
24:  $\mathit{ncss}[K]\leftarrow[\,]$ $\triangleright$ Clear size $K$ nested combinations
25:end for
26:return $t_{opt}$
@@
The main difference between the brute-force specification 5 and Algorithm 3 lies in the order of execution. In the specification, functions are composed sequentially, meaning each function begins only after the previous one has completed. If written as pseudo-code, both $\mathit{min}{}_{E_{\text{0-1}}}$ and $\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}$ would appear outside the outermost loop that defines $\mathit{\mathit{nestedCombs}}\left(K,G\right)$. This is inefficient because $\mathit{\mathit{nestedCombs}}\left(K,G\right)$ generates a prohibitively large number of candidate solutions.
@@ 4.1 Coreset selection method
Algorithm 3 Coreset Selection ($\mathit{hodtCoreset}(K,M,\mathit{xs},\mathit{BS},R,L,B_{\max},c)$)
1:Input: Coreset parameters: $\mathit{BS}$: block size; $R$: shuffle iterations; $L$: max-heap size; $B_{\max}$: max input size; $c\in(0,1]$: shrinking factor
2:      $\mathit{hodt}$ parameters: $K$: splitting rules; $M$: polynomial degree; $\mathit{xs}$: data list
3:Output: Max-heap with top $L$ configurations and data blocks
[…]
5:while $|\mathcal{C}|\leq B_{\max}$ do
6:  $\mathcal{C}_{B}\leftarrow\{C_{1},C_{2},\ldots,C_{\lceil|\mathcal{C}|/\mathit{BS}\rceil}\}$ $\triangleright$ Reshuffle and divide into blocks
7:  $\mathcal{H}_{L}\leftarrow\emptyset_{L}$ $\triangleright$ Initialize size $L$ max-heap
8:  for $r\in\mathit{range}(1,R)$ do
9:   for $C\in\mathcal{C}_{B}$ do
10:    $\mathit{cnfg}\leftarrow\mathit{hodt}(K,M,C)$ $\triangleright$ Compute configuration
11:    $\mathcal{H}_{L}.\mathit{push}(\mathit{cnfg},C)$ $\triangleright$ Push to heap
12:   end for
13:  end for
14:  $\mathcal{C}\leftarrow\mathit{unique}(\mathcal{H}_{L})$ $\triangleright$ Merge blocks, remove duplicates
15:  $L\leftarrow L\times c$ $\triangleright$ Shrink heap size
16:end while
[…]
19:return $\mathcal{H}_{L}$
@@ 4.2 $\mathit{sodt}$ with selected hyperplanes
Algorithm 4 $\mathit{sodt}$ with selected hyperplanes ($\mathit{sodtWSH}(K,M,\alpha,\mathit{xs})$)
1:Input: $K$: splitting rules; $M$: polynomial degree; $\alpha$: duplicate threshold; $\mathit{xs}$: data list
2:Output: Nested combinations $\mathit{NCs}$ without crossed hyperplanes, as $(k+2)\times(k+1)$ matrices
[…]
7:for $n\in\mathit{range}(0,N)$ do
8:  for $k\in\mathit{reverse}(\mathit{range}(K,n+1))$ do
9:   $\mathit{ncss}^{\prime}\leftarrow\mathit{filter}_{q(\alpha)}(\mathit{map}(\cup[n],\mathit{ncss}[k-1]))$ $\triangleright$ Filter duplicates below $\alpha$
10:   $t_{best}[k]\leftarrow\min_{E_{0-1}}(\mathit{mapL}(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}^{\prime}))$ $\triangleright$ Compute best configuration
11:   $\mathit{ncss}[k]\leftarrow\mathit{ncss}[k]\cup\mathit{reverse}(\mathit{ncss}^{\prime})$ $\triangleright$ Update combinations
12:   $\mathit{ncss}[K]\leftarrow[\,]$ $\triangleright$ Clear size $K$ combinations
13:  end for
14:end for
15:return $t_{best}$
@@ 5.1 Computational analysis
In this section, we analyze in detail the computational efficiency of $\mathit{sodt}$ for solving the size-constrained ODT problem. Since $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ share similar advantages—both are fully vectorized but unable to exploit efficient dynamic programming (DP)—we focus on comparing $\mathit{sodt}_{\text{vec}}$ against the dynamic programming algorithm $\mathit{sodt}_{\text{rec}}$.
@@ 5.1.1 Comparison between the vectorized and recursive implementation
Figure 5: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on sequential setting.
Figure 6: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on parallel setting.
@@
Following the detailed comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ in both sequential and parallel settings, we now turn to the computational scalability of the winning method, $\mathit{sodt}_{\text{vec}}$. We focus on analyzing the pure computational scalability of the algorithm by benchmarking its performance without any acceleration techniques. Specifically, we measure the total wall-clock runtime of $\mathit{sodt}_{\text{vec}}$ for solving $1\times 10^{6}$ instances of feasible nested combinations (with no crossed hyperplanes) for different values of $K$. In our setup, cases with $K\leq 4$ are executed on the CPU, while cases with $K>4$ are executed on the GPU.
@@ 5.2 Combinatorial complexity of the hyperplane/hypersurface decision tree
5.2.1 Combinatorial complexity of nested combinations after filtered out crossed hyperplanes
Figure 8: Combinatorial complexity of feasible nested combinations with varying $K$. Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) to the value of $K$ labeled below each group of bars.
@@ 5.2.1 Combinatorial complexity of nested combinations after filtered out crossed hyperplanes
Comparing the combinatorial complexity of hypersurface and hyperplanes in the same dimension
Figure 9: Combinatorial complexity of feasible nested combinations with varying $K$. Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) to the value of $K$ labeled below each group of bars.
@@ 5.3 Computational experiments with synthetic datasets
2 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.12 (0.04) 93.82 (3.01) 91.72 (3.83) 100 (0.00) 100 (0.00) 82.61 (0.04) 86.12 (3.73) 84.57 (4.03) 93.50 (3.21) 100 (0.00) 2.92 (0.27) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.49 (0.16) 2.00 (0.00) 1.67 (1.20) 1.67 (1.20) 1.96 (0.15) 2.15 (0.20) 2.00 (0.00) 1.65 (0.16) 1.66 (0.16)
3 CART-depth-3 CART-size-6 ConTree-3 HODT Ground Truth 90.72 (047) 93.88 (3.89) 95.98 (2.93) 100 (0.00) 100 (0.00) 81.30 (5.71) 83.92 (5.55) 85.25 (4.96) 94.91 (2.36) 100 (0.00) 5.90 (0.00) 6.00 (0.00) 6.94 (0.24) 3.00 (0.00) 3.00 (0.00) 3.00 (0.00) 4.02 (0.65) 3.00 (0.00) 3.00 (0.00) 2.72 (0.45) 2.83 (0.14) 3.10 (0.24) 2.99 (0.03) 2.25 (0.00) 2.18 (0.11) 2.77 (0.19) 2.71 (0.20) 2.99 (0.04) 2.18 (0.25) 2.07 (0.24)
4 CART-depth-3 CART-size-8 ConTree-3 HODT Ground Truth 87.98 (5.26) 93.90 (3.93) 93.12 (4.02) 97.46 (1.70) 100 (0.00) 77.66 (5.38) 81.93 (4.93) 81.51 (4.78) 89.89 (3.71) 100 (0.00) 6.12 (90.86) 8.00 (0.00) 7.00 (0.00) 4.00 (0.00) 4.00 (0.00) 3.00 (0.00) 4.54 (0.57) 3.00 (0.00) 3.52 (0.50) 3.52 (0.50) 2.87 (0.14) 3.49 (0.21) 3.00 (0.00) 2.63 (0.19) 2.63 (0.19) 2.86 (0.15) 3.13 (0.21) 3.00 (0.00) 2.60 (0.31) 2.63 (0.33)
@@
100 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.76 (5.11) 97.62 (2.51) 93.52 (3.53) 100 (0.00) 100 (0.00) 75.46 (15.70) 80.26 (16.80) 78.62 (16.14) 84.38 (19.27) 100 (0.00) 2.88 (0.32) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 0.00 (0.00) 2.00 (0.00) 3.92 (0.69) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.96 (0.11) 3.01 (0.30) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.96 (0.12) 2.52 (0.33) 2.00 (0.00) 1.66 (0.25) 1.70 (0.25)
200 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.00 (4.49) 94.49 (3.02) 92.87 (3.38) 99.94 (0.16) 100 (0.00) 84.62 (5.23) 89.57 (4.04) 88.51 (3.99) 97.97 (1.01) 100 (0.00) 2.92 (0.27) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.26 (0.44) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.50 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.98 (0.07) 2.17 (0.21) 2.00 (0.00) 1.71 (0.25) 1.70 (0.25)
400 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 87.45 (4.69) 92.58 (3.48) 91.44 (3.37) 99.85 (0.21) 100 (0.00) 84.62 (4.90) 89.35 (4.08) 88.78 (4.00) 98.98 (0.58) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 2.53 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.99 (0.05) 2.15 (0.24) 2.00 (0.00) 1.71 (0.25) 1.70 (0.25)
800 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 86.91 (4.63) 92.21 (3.43) 91.00 (3.59) 99.87 (0.14) 100 (0.00) 85.14 (4.96) 90.12 (3.82) 89.18 (3.94) 99.40 (0.36) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.3 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.53 (0.18) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.13 (0.22) 2.00 (0.00) 1.69 (0.25) 1.70 (0.25)
1600 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 85.66 (5.05) 91.48 (3.69) 90.13 (3.69) 99.87 (0.12) 100 (0.00) 84.97 (5.15) 90.53 (3.71) 89.38 (3.84) 99.65 (0.22) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.51 (0.16) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.16 (0.26) 2.00 (0.00) 1.70 (0.24) 1.70 (0.25)